Grade 11Mechanics

the resultant of 2 forces is 2p and √2p then the angle between the force

shivam
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
let a and b be the two forces,
then
a + b = 2p (let)
a – b  = √2p (let)
solving both resultants we get,
2a = p(2 + √2)
a =  p(2 + √2)/2
and 
b =  p(2 - √2)/2
cos@ = a.b/|a||b| =   p(2 + √2)/2. p(2 - √2)/2)/2*|p|^2